# Graph sensitivity under join and decomposition

### v1

@@ 3.2. Joining an empty graph to a nonempty graph
Since $\sigma(G_{n})\to\infty$ as $n\to\infty$, we observe that $\alpha(G_{n})\to\infty$ as $n\to\infty$. So there exists an $N_{1}>0$ such that for any $n\geq N_{1}$ we have $\alpha(G_{n})\geq m$. This means that for each $n\geq N_{1}$, $\sigma(\overline{K}_{m}\vee G_{n})$ equals either $\left\lceil\frac{\alpha(G_{n})+1}{2}\right\rceil$, $\alpha(G_{n})+1-m$, or $\sigma(G)$ by Theorem 3.7 (3).
Now let $C>0$. Each of the quantities $\left\lceil\frac{\alpha(G_{n})+1}{2}\right\rceil$, $\alpha(G_{n})+1-m$, and $\sigma(G_{n})$ become arbitrarily large as $n$ increases. Thus, there exists an $N_{2}\geq N_{1}$ such that for any $n\geq N_{2}$, each of these three quantities is greater than $C$. Thus, for any $n\geq N_{2}$, $\sigma(\overline{K}_{m}\vee G_{n})\geq C$, which completes the proof. ∎
@@ 3.3. Joining an empty graph to a regular nonempty graph
If $m$ is large relative to the number of vertices of a graph $G$, then $\sigma(\overline{K}_{m}\vee G)$ is given by Theorem 3.7. The following proposition shows that when $G$ is regular and $m$ is sufficiently large, we can obtain a simple formula for $\sigma(\overline{K}_{m}\vee G)$.
@@ 4. Sensitivity and stable blocks
In this section we consider when a graph $G$ has a decomposition into sets called stable blocks. These are defined by Larson in [21] and allow isolating maximum independent sets and describing sensitivity.
@@
Note that if $S$ is a stable block, then $\mathrm{int}(S)$ is nonempty. Stable blocks were used to efficiently reduce independence number calculations to subgraphs [20]. We use them to bound sensitivity in Section 4.1, and in Theorem 4.19 to compute sensitivity precisely in terms of the blocks when there is a full partition of $V_{G}$ into such sets. We also define vertex identification and use it as a means to construct a stable block in Corollary 4.6, repeat this to bound the sensitivity of a rooted product in Section 4.2, and indicate how to construct a graph with a partition into stable blocks in Section 4.3. As an application we compute the sensitivity of a corona of two graphs and of families of successive corona graphs.
@@ 4.1. Single stable block
For the star graph $S_{n}=K_{1}\vee\overline{K}_{n}$ with $\{v\}$ the distinguished vertex in $K_{1}$, and the cycle graph $C_{m}$ with $m\geq 3$ and any distinguished vertex $\{u\}$, the cycle-star graph [7] is $CS_{m,n}=(C_{m}+S_{n})/\{u,v\}$. It has $\alpha(CS_{m,n})=\left\lfloor\frac{m}{2}\right\rfloor+n$ and $\Delta(CS_{m,n})=n+2$. Since there is a maximum independent subset in $C_{m}$ not containing $u$ and $CS_{m,n}[\mathrm{int}(C_{m})]=C_{m}\setminus\{u\}$ is a nonempty path, we conclude that $\sigma(CS_{m,n})=\sigma(CS_{m,n}[\mathrm{int}(C_{m})])=1$ for $m\geq 5$ (so that the path has length at least $4$) and the family $\{CS_{m,n}\}_{n=1}^{\infty}$ is insensitive.
@@
The following definition used in [20] was originally given in [31].
@@
It is possible that the empty set is the only maximum critical independent set, as is the case when $G=K_{n}$ for all $n\geq 3$. But if not, the Independence Decomposition Theorem of [20] produces a stable block.
@@
At the start of the proof of Theorem $2.4$ in [20], Larson proves that $I_{c}$ is a maximum independent subset of $G[X]$, so $\alpha(G[X])=|I_{c}|$.
@@
Larson also proves that the set $X$ is independent of the choice of maximum critical independent set [20] and that a maximum critical independent set can be found in polynomial time [19]. Hence when there exists a nonempty maximum critical independent set Proposition 4.11 provides an effective way to produce a stable block in a graph and use Proposition 4.2 to obtain an upper bound on sensitivity, which is most useful for proving a family is insensitive.
@@ 4.2. Rooted products and coronas
Given a graph $G$ of order $n$ and a set of graphs $\mathcal{H}=\{H_{1},\dots,H_{n}\}$ rooted at $v_{1},\dots,v_{n}$ respectively, each root can be identified with a vertex of $G$ to obtain the rooted product $G(H_{1},\dots,H_{n})$ of $\mathcal{H}$ by $G$. This construction was originally introduced in [12] where the characteristic polynomial of its adjacency matrix was determined in terms of the characteristic polynomials for $H_{1},\dots,H_{n}$.
@@
For any graph $G$ with one or more distinguished vertices, using any $1$-cone $H$ whose interior is nonempty results in a wide variety of examples by applying Corollary 4.14 at the chosen vertices of $G$. For instance, if $H$ is the join of $K_{1}$ and $K_{n}$, $P_{n}$, or $C_{n}$ for $n\geq 2$, $n\geq 4$, or $n\geq 5$ respectively, as in Example 2.3, then the resulting graph will have sensitivity $1$, as will any tree built by successive vertex identification involving at least one path of length at least $4$. In particular the royal petunia graphs of  [22] contain a copy of $K_{1}\vee P_{n}$ as a stable block, and thus have sensitivity $1$.
If we identify a copy of $K_{1}\vee H$ rooted at the vertex in $K_{1}$ to each vertex in $G$ then the resulting rooted product is called the corona $G\odot H$ of $G$ and $H$. The corona operation was introduced in [10] to produce a graph whose automorphism group could be easily described as a wreath product involving the automorphism groups of the component graphs.
@@
If $H$ is nonempty then $G\odot H=G(K_{1}\vee H)$ will satisfy the hypotheses of Corollary 4.14 and each copy of $K_{1}\vee H$ will be a stable block in $G\odot H$, so $\sigma(G\odot H)\leq\sigma(K_{1}\vee H)=\sigma(H)$. For example, the Type III windmill graphs in [18] are $K_{m}\odot K_{n}$ and hence by Corollary 4.14 have sensitivity $\sigma(K_{m}\odot K_{n})=1$. Since for $m\geq 2$ we have
@@ 5. Concluding remarks
In a generalized join $G[H_{1},\dots,H_{n}]$, each subgraph induced by $H_{i}$, $1\leq i\leq n$ is a module, i.e., a set of vertices that have the same neighbors outside $H_{i}$. Thus $H_{1},\dots,H_{n}$ forms a modular partition of the full graph, and conversely any modular partition of a graph can be used to express it as a generalized join. In general it is possible to partition the vertices of a graph into modules in multiple ways. The modular decomposition tree of a graph $G$ exhibits the nesting of its strong modules (see [14]), uniquely determines $G$ (see [3]), and the maximal proper modules in that tree form a partition of the vertex set of $G$ that is referred to as a maximal modular partition of $G$ (see [15]). See [3] for definitions, a list of several other terms that have been used for similar ideas in a variety of contexts, and precise statements of a modular decomposition theorem and related results. For a brief primer and an interesting new generalization see [14] and for further details and a reference list see [15].
When $G$ has more than one connected component $H_{1},\dots,H_{n}$ the unique maximal modular partition is the parallel decomposition $G=\overline{K}_{n}[H_{1},\dots,H_{n}]$, for which we described the sensitivity in Example 3.14. When the complement $G^{\prime}$ has more than one connected component $H_{1},\dots,H_{n}$ then $G=K_{n}[H_{1},\dots,H_{n}]$ is the series decomposition, for which we described sensitivity in the lexicographic product case, $K_{n}\circ H$, in Example 3.15. When $G$ and $G^{\prime}$ are connected, then for a maximal modular partition $H_{1},\dots,H_{n}$ of $G$ we can replace each $H_{i}$ with a single vertex and obtain a quotient $Q$ that has no nontrivial modules and allows us to write $G$ as a generalized join $G=Q[H_{1},\dots,H_{n}]$. For example, in the parallel and series decompositions the graphs $\overline{K}_{n}$ and $K_{n}$ are quotients of $G$ by $H_{1},\dots,H_{n}$. This suggests limitations on finding the sensitivity of a generalized join explicitly in terms of its components, since in full generality this would amount to finding the sensitivity of a general graph.
@@ 6. Acknowledgements
We thank Davis Bolt and Gabriel Ory for computing support carried out in SageMath [29]. The Idaho State University Career Path Internship Program provided funding for that computing support and some work by Jacob Tolman.

### v2

@@ 3.2. Joining an empty graph to a nonempty graph
Since $\sigma(G_{n})\to\infty$ as $n\to\infty$, we observe that $\alpha(G_{n})\to\infty$ as $n\to\infty$. So there exists an $N_{1}>0$ such that for any $n\geq N_{1}$ we have $\alpha(G_{n})\geq m$. This means that for each $n\geq N_{1}$, $\sigma(\overline{K}_{m}\vee G_{n})$ equals either $\left\lceil\frac{\alpha(G_{n})+1}{2}\right\rceil$, $\alpha(G_{n})+1-m$, or $\sigma(G_{n})$ by Theorem 3.7 (3).
Now let $C>0$. Each of the quantities $\left\lceil\frac{\alpha(G_{n})+1}{2}\right\rceil$, $\alpha(G_{n})+1-m$, and $\sigma(G_{n})$ become arbitrarily large as $n$ increases. Thus, there exists an $N_{2}\geq N_{1}$ such that for any $n\geq N_{2}$, each of these three quantities is greater than $C$ and hence $\sigma(\overline{K}_{m}\vee G_{n})\geq C$, which completes the proof. ∎
@@ 3.3. Joining an empty graph to a regular nonempty graph
If $m$ is large relative to the number of vertices of a graph $G$, then $\sigma(\overline{K}_{m}\vee G)$ is given by Theorem 3.7. The following proposition shows that when $G$ is regular and $m$ is sufficiently large, there is a simple formula for $\sigma(\overline{K}_{m}\vee G)$.
@@ 4. Sensitivity and stable blocks
In this section we consider when a graph $G$ has a decomposition into sets called stable blocks. These are defined by Larson in [Larson-12] and allow isolating maximum independent sets and describing sensitivity.
@@
Note that if $S$ is a stable block, then $\mathrm{int}(S)$ is nonempty. Stable blocks were used to efficiently reduce independence number calculations to subgraphs [Larson-11]. We use them to bound sensitivity in Section 4.1, and in Theorem 4.19 to compute sensitivity precisely in terms of the blocks when there is a full partition of $V_{G}$ into such sets. We also define vertex identification and use it as a means to construct a stable block in Corollary 4.6, repeat this to bound the sensitivity of a rooted product in Section 4.2, and indicate how to construct a graph with a partition into stable blocks in Section 4.3. As an application we compute the sensitivity of a corona of two graphs and of families of successive corona graphs.
@@ 4.1. Single stable block
For the star graph $S_{n}=K_{1}\vee\overline{K}_{n}$ with $\{v\}$ the distinguished vertex in $K_{1}$, and the cycle graph $C_{m}$ with $m\geq 3$ and any distinguished vertex $\{u\}$, the cycle-star graph [Clark-85] is $CS_{m,n}=(C_{m}+S_{n})/\{u,v\}$. It has $\alpha(CS_{m,n})=\left\lfloor\frac{m}{2}\right\rfloor+n$ and $\Delta(CS_{m,n})=n+2$. Since there is a maximum independent subset in $C_{m}$ not containing $u$ and $CS_{m,n}[\mathrm{int}(C_{m})]=C_{m}\setminus\{u\}$ is a nonempty path, we conclude that $\sigma(CS_{m,n})=\sigma(CS_{m,n}[\mathrm{int}(C_{m})])=1$ for $m\geq 5$ (so that the path has length at least $4$) and the family $\{CS_{m,n}\}_{n=1}^{\infty}$ is insensitive.
@@
The following definition used in [Larson-11] was originally given in [Zhang-90].
@@
It is possible that the empty set is the only maximum critical independent set, as is the case when $G=K_{n}$ for all $n\geq 3$. But if not, the Independence Decomposition Theorem of [Larson-11] produces a stable block.
@@
At the start of the proof of Theorem $2.4$ in [Larson-11], Larson proves that $I_{c}$ is a maximum independent subset of $G[X]$, so $\alpha(G[X])=|I_{c}|$.
@@
Larson also proves that the set $X$ is independent of the choice of maximum critical independent set [Larson-11] and that a maximum critical independent set can be found in polynomial time [Larson-07]. Hence when there exists a nonempty maximum critical independent set Proposition 4.11 provides an effective way to produce a stable block in a graph and use Proposition 4.2 to obtain an upper bound on sensitivity, which is most useful for proving a family is insensitive.
@@ 4.2. Rooted products and coronas
Given a graph $G$ of order $n$ and a set of graphs $\mathcal{H}=\{H_{1},\dots,H_{n}\}$ rooted at $v_{1},\dots,v_{n}$ respectively, each root can be identified with a vertex of $G$ to obtain the rooted product $G(H_{1},\dots,H_{n})$ of $\mathcal{H}$ by $G$. This construction was originally introduced in [Godsil-McKay-78] where the characteristic polynomial of its adjacency matrix was determined in terms of the characteristic polynomials for $H_{1},\dots,H_{n}$.
@@
For any graph $G$ with one or more distinguished vertices, using any $1$-cone $H$ whose interior is nonempty results in a wide variety of examples by applying Corollary 4.14 at the chosen vertices of $G$. For instance, if $H$ is the join of $K_{1}$ and $K_{n}$, $P_{n}$, or $C_{n}$ for $n\geq 2$, $n\geq 4$, or $n\geq 5$ respectively, as in Example 2.3, then the resulting graph will have sensitivity $1$, as will any tree built by successive vertex identification involving at least one path of length at least $4$. In particular the royal petunia graphs of  [Latyshev-Kokhas-23] contain a copy of $K_{1}\vee P_{n}$ as a stable block, and thus have sensitivity $1$.
If we identify a copy of $K_{1}\vee H$ rooted at the vertex in $K_{1}$ to each vertex in $G$ then the resulting rooted product is called the corona $G\odot H$ of $G$ and $H$. The corona operation was introduced in [Frucht-Harary-70] to produce a graph whose automorphism group could be easily described as a wreath product involving the automorphism groups of the component graphs.
@@
If $H$ is nonempty then $G\odot H=G(K_{1}\vee H)$ will satisfy the hypotheses of Corollary 4.14 and each copy of $K_{1}\vee H$ will be a stable block in $G\odot H$, so $\sigma(G\odot H)\leq\sigma(K_{1}\vee H)=\sigma(H)$. For example, the Type III windmill graphs in [Kooij-19] are $K_{m}\odot K_{n}$ and hence by Corollary 4.14 have sensitivity $\sigma(K_{m}\odot K_{n})=1$. Since for $m\geq 2$ we have
@@ 5. Concluding remarks
In a generalized join $G[H_{1},\dots,H_{n}]$, each subgraph induced by $H_{i}$, $1\leq i\leq n$ is a module, i.e., a set of vertices that have the same neighbors outside $H_{i}$. Thus $H_{1},\dots,H_{n}$ forms a modular partition of the full graph, and conversely any modular partition of a graph can be used to express it as a generalized join. In general it is possible to partition the vertices of a graph into modules in multiple ways. The modular decomposition tree of a graph $G$ exhibits the nesting of its strong modules (see [HMSZ-24]), uniquely determines $G$ (see [Brandstadt-Le-Spinrad-99]), and the maximal proper modules in that tree form a partition of the vertex set of $G$ that is referred to as a maximal modular partition of $G$ (see [Habib-Paul-10]). See [Brandstadt-Le-Spinrad-99] for definitions, a list of several other terms that have been used for similar ideas in a variety of contexts, and precise statements of a modular decomposition theorem and related results. For a brief primer and an interesting new generalization see [HMSZ-24] and for further details and a reference list see [Habib-Paul-10].
When $G$ has more than one connected component $H_{1},\dots,H_{n}$ the unique maximal modular partition is the parallel decomposition $G=\overline{K}_{n}[H_{1},\dots,H_{n}]$, for which we described the sensitivity in Example 3.14. When the complement $\overline{G}$ has more than one connected component $H_{1},\dots,H_{n}$ then $G=K_{n}[H_{1},\dots,H_{n}]$ is the series decomposition, for which we described sensitivity in the lexicographic product case, $K_{n}\circ H$, in Example 3.15. When $G$ and $\overline{G}$ are connected, then for a maximal modular partition $H_{1},\dots,H_{n}$ of $G$ we can replace each $H_{i}$ with a single vertex and obtain a quotient $Q$ that has no nontrivial modules and allows us to write $G$ as a generalized join $G=Q[H_{1},\dots,H_{n}]$. For example, in the parallel and series decompositions the graphs $\overline{K}_{n}$ and $K_{n}$ are quotients of $G$ by $H_{1},\dots,H_{n}$. This suggests limitations on finding the sensitivity of a generalized join explicitly in terms of its components, since in full generality this would amount to finding the sensitivity of a general graph.
@@ 6. Acknowledgements
We thank Davis Bolt and Gabriel Ory for computing support carried out in SageMath [sagemath]. The Idaho State University Career Path Internship Program provided funding for that computing support and some work by Jacob Tolman.